At a time t seconds after it is thrown up in the air, a tomato is
at a height (in meters) of f(t) = =4.9t
2 +50t +5 m.
A. What is the average velocity of the tomato during the
first 4 seconds? (Include units.)
B. Find (exactly) the instantaneous velocity of the tomato
at t = 4. (Include units.)
C. What is the acceleration at t = 4? (Include units.)
D. How high does the tomato go? (Include units.)
E. How long is the tomato in the air?
1
Expert's answer
2020-03-13T10:16:09-0400

A. "\bar{V}= \frac{f(4)-f(0)}{4}=30.4 \frac{m}{c}"

B. "V(t)=f_t'=-9.8t+50" hence "V(4)=10.8\frac{m}{s}"

C. "a(t)=V_t'=-9.8\frac{m}{s^2}"

D. If"\tau" is time up then "V(\tau)=0" hence "-9.8\tau+50=0" then "\tau=5.1s" then "f(\tau)=H\approx132.6m"

E. If "t" is time down then "H=\frac{gt^2}{2}" hence "t=\sqrt \frac{2H}{g}\approx5.1s" then "T=t+\tau\approx10.2s"